# I Dream My Painting: Connecting MLLMs and Diffusion Models via Prompt Generation for Text-Guided Multi-Mask Inpainting

### v1

@@ 1 Introduction
This paper proposes a new pipeline to address the mentioned challenges. First, we introduce a new task called text-guided multi-mask inpainting, which involves inpainting multiple image regions simultaneously, each guided by a specific text prompt. We go further by exploring the automatic generation of textual prompts directly from a corrupted image, as done in Brush2Prompt [6], but with the added complexity of predicting multiple prompts for different regions in a single pass. We frame this as a visual instruction-following task for multimodal large language models (MLLMs) like LLaVA [21]. By fine-tuning this model using QLoRA [10], we automatically generate multi-prompts for use as user suggestions or direct text prompts for Stable Diffusion. Finally, we adapt the rectified cross-attention technique from FreestyleNet [38] and concatenate the multi-mask prompts into a single prompt, enforcing the correct parts of the prompt on their corresponding regions in the diffusion model.
We focus our analysis on a dataset of digitized paintings from WikiArt, as these present a challenging application. Paintings often feature complex imagery and abstract concepts, demanding creativity for interpretation and identification, complicating prompt generation and text-guided inpainting. Since these images lack object-level annotations, we demonstrate that such annotations can be obtained using MLLMs like Kosmos-2 [29] and LLaVA, showcasing our pipeline’s applicability to datasets without annotations. To illustrate the method’s adaptability to other domains, such as photographic scenes, we also test it on the Densely Captioned Images dataset [35], which includes region-level annotations.
@@ 4.2 Results and Discussion
FID ↓ LPIPS ↓ PSNR ↑ CLIP-IQA ↑ CLIPSim-I2I ↑ CLIPSim-T2I ↑
SD-2-Inp-HQPrompt 19.18 (31.86) 22.82 (24.29) 14.55 (14.36) 71.51 (73.89) 84.87 (85.63) 21.10 (20.92)
[…]
Table 2: Inpainting results. We compared our solution (SD-2-Inp-RCA-FineTuned), which leverages Stable Diffusion 2.0-Inpainting fine-tuned with RCA, against various non-fine-tuned and fine-tuned approaches. Results are reported for the full test set and the subset of test examples with more than one mask (multi-mask, shown in parentheses).
Data Annotation Quality
Prompt Generation Quality and Plausibility
We report prompt generation results in Table 1. We compared our fine-tuning approach for LLaVA, which includes mask coloring and prompt-answer formatting, against the base LLaVA model’s visual instruction-following abilities. The base model struggles with multi-mask prediction, often avoiding the question and providing a simple image description instead. To simplify the task, we evaluated the base model in a single-mask setting by masking a random object and using the same prompt from the Brush2Prompt paper (LLaVA-Prompt). Even in this scenario, the model inaccurately describes hidden objects, focusing on visible entities or the entire image, leading to lower scores than our fine-tuned versions.
We fine-tuned two LLaVA variants with our method: LLaVA-1Mask, which randomly selects one entity and uses a single gray mask per training example, and LLaVA-MultiMask, which is trained in a multi-mask setting. We assessed LLaVA-1Mask in the single-mask prompt generation scenario and evaluated LLaVA-MultiMask for single-mask prediction (LLaVA-MultiMask-1Pred), selecting the same object as LLaVA-1Mask for each test example. Surprisingly, the multi-mask model outperforms even in the single-mask setting, suggesting that training on the more complex multi-mask task enhances prompt generation.
We also compared LLaVA-MultiMask by selecting all objects in an image for the test and placing the same test object from the previous alternatives as the last object in the prompt. We then simultaneously generated prompts for all masks, evaluating only the last mask (LLaVA-MultiMask-LastPred). This experiment shows minimal performance degradation for the multi-mask model despite having access to approximately 18% fewer unmasked regions on average.
On DCI (Table 4), the prompts generated by the base model achieve CLIPSim results comparable to LLaVA-MultiMask but fall significantly short in accuracy and BLEU. We attribute the similar CLIPSim scores to DCI’s annotations, which are longer, often truncated by the CLIP text encoder, and more complex than the base model’s predictions, lowering the CLIPSim score.
Temperature and Diversity
Figure 5: Qualitative results. We compare the different approaches tested in our evaluation. Fine-tuning with RCA enables the model to better follow prompts, capture details, and respect inpainting locations while reducing susceptibility to concept blending. Using generated prompts with Stable Diffusion produces creative and plausible results, in contrast to the alternative without prompts, which tends to fill areas with background textures or unusual objects from the model’s initial distribution. The second example is from DCI.
@@
We provide two baselines where we tried SD-2-Inpainting as-is with the general prompt “High quality” (SD-2-Inp-HQPrompt) or with the multi-mask prompt (SD-2-Inp). We notice that the second model achieves the best FID, LPIPS, and PSNR scores at the cost of not closely adhering to the prompts, as shown by the CLIPSim scores. After further inspection, we noticed that the SD-2-Inp model tends to blend all the concepts in a single prompt, sometimes filling specific masks with the background or unrecognizable objects. Thus, evaluating only the first three metrics is unsuitable for this task.
@@
We evaluated the entire pipeline using our fine-tuned Stable Diffusion with RCA and generated prompts. As expected, this model achieves the second-best FID score and outperforms SD-2-Inp-HQPrompt in inpainting ground truth entities, showing significant improvement even without access to ground truth prompts.
The improved region-prompt alignment from RCA is confirmed by results on DCI (Table 4), where the RCA variant of the model outperforms the alternatives in text-to-image CLIPSim.
@@ A Additional Dataset Information
This section provides additional details on the automatically annotated dataset of digitized images of paintings from $\mathcal{A}rt\mathcal{G}raph$. As described in the main paper, the images were downloaded as thumbnails from the WikiArt API, using the formats HalfHD, Large, and PinterestLarge. For each image, we selected the highest resolution available. The counts and resolutions of the images are summarized in Table 5.
Figure 6: Comparative analysis on multi-mask inpainting.
@@ B.1 Analysis on the Number of Masks
We compare the base Stable Diffusion-2-Inpainting model and our RCA fine-tuned version, using generated prompts to evaluate performance across varying numbers of masks (Fig. 6). The figure demonstrates that our method consistently achieves superior region-prompt alignment compared to the base Stable Diffusion model, regardless of the number of masks. Notably, at $N=5$, the curves begin to converge, suggesting that multi-mask inpainting is particularly effective for cases with fewer masks ($N\geq 1$). This could be because as the number of masks increases, the inpainting area expands, leading to more intersections between masks and longer prompts containing multiple sub-prompts. This complexity may reduce the precision of metric calculations, potentially diminishing the measurable advantages of RCA.

### v2

@@ 1 Introduction
This paper proposes a new pipeline to address the mentioned challenges. First, we introduce a new task called text-guided multi-mask inpainting, which involves inpainting multiple image regions simultaneously, each guided by a specific text prompt. We go further by exploring the automatic generation of textual prompts directly from a corrupted image, as done in Brush2Prompt [6], but with the added complexity of predicting multiple prompts for different regions in a single pass. We frame this as a visual instruction-following task for a multimodal large language model (MLLM) like LLaVA [21]. By fine-tuning this model using QLoRA [10], we automatically generate multiple prompts for use as user suggestions or direct text prompts for Stable Diffusion. Finally, we adapt the rectified cross-attention technique from FreestyleNet [38] and concatenate the multi-mask prompts into a single prompt, enforcing the correct parts of the prompt on their corresponding regions in the diffusion model, achieving similar quantitative results and more aesthetically pleasing qualitative outcomes compared to inpainting masks individually.
We focus our analysis on a dataset of digitized artworks from WikiArt, as these present a challenging application. Artworks often feature complex imagery and abstract concepts, demanding creativity for interpretation and identification, complicating prompt generation and text-guided inpainting. Since these images lack object-level annotations, we demonstrate that such annotations can be obtained using MLLMs like Kosmos-2 [29] and LLaVA, showcasing our pipeline’s applicability to datasets without annotations. To illustrate the method’s adaptability to other domains, such as photographic scenes, we also test it on the Densely Captioned Images dataset [35], which includes region-level annotations.
@@ 4.1.2 Implementation Details
FID ↓ LPIPS ↓ PSNR ↑ CLIP-IQA ↑ CLIPSim-I2I ↑ CLIPSim-T2I ↑
[…]
SD-2-Inp-FineTuned 15.49 (27.83) 22.06 (23.96) 14.44 (14.06) 74.64 (77.68) 89.05 (89.04) 26.31 (25.40)
[…]
SD-2-Inp-Repeated-FineTuned 15.79 (29.20) 22.73 (25.09) 14.43 (14.11) 72.29 (73.12) 88.85 (88.91) 26.95 (26.23)
[…]
SD-2-Inp-RCA-FineTuned-Gen 15.30 (27.94) 22.69 (24.42) 14.05 (13.64) 72.80 (76.01) 87.47 (87.68) 23.25 (22.94)
Table 3: Inpainting results. We compared our solution (SD-2-Inp-RCA-FineTuned), which leverages Stable Diffusion 2.0-Inpainting fine-tuned with RCA, against various non-fine-tuned and fine-tuned approaches. Results are reported for the full test set and for test examples with multiple masks (in parentheses). The gray rows show alternatives that repeat the inpainting process for each mask.
4.1.3 Evaluation Metrics
4.2 Results and Discussion
Data Annotation Quality
Prompt Generation Quality and Plausibility
Table 1 compares our fine-tuning approach for LLaVA, which includes mask coloring and prompt-answer formatting, against the base LLaVA model’s visual instruction-following abilities. The base model, limited to visual instruction-following, struggles with multi-mask prediction, often defaulting to general image descriptions. Even in a simplified single-mask scenario (using the Brush2Prompt paper’s prompt), it inaccurately describes hidden objects, focusing on visible entities, leading to lower scores than our fine-tuned variants.
We fine-tuned two models: LLaVA-1Mask, trained on single-mask examples, and LLaVA-MultiMask, trained in a multi-mask setting. LLaVA-1Mask performs well in single-mask predictions, but LLaVA-MultiMask surprisingly outperforms it, even in single-mask scenarios (1Pred), suggesting that multi-mask training enhances prompt generation capabilities. Additionally, we evaluated LLaVA-MultiMask in a full multi-mask test, focusing on a single selected mask (LLaVA-MultiMask-LastPred). Despite having access to 18% fewer unmasked regions on average, performance degradation was minimal, underscoring the robustness of the multi-mask model.
On DCI (Table 2), while the base model’s prompts achieve CLIPSim scores comparable to LLaVA-MultiMask, they fall short in accuracy and BLEU. This likely stems from DCI’s complex annotations, which are often truncated by the CLIP encoder, reducing CLIPSim for both models.
Temperature and Diversity
Figure 5: Qualitative results. We compare the different approaches tested in our evaluation. Fine-tuning with RCA enables the model to better follow prompts, capture details, and respect inpainting locations while reducing susceptibility to concept blending. The second example is from DCI.
@@ 4.2 Results and Discussion
We provide two baselines where we tried SD-2-Inpainting as-is with the general prompt “High quality” (SD-2-Inp-HQPrompt) or with the multi-mask prompt (SD-2-Inp). We notice that the second model achieves the best FID, LPIPS, and PSNR at the cost of not closely adhering to the prompts, as shown by the CLIPSim scores. After further inspection, we noticed that the SD-2-Inp model tends to blend all the concepts in a single prompt, sometimes filling specific masks with the background or unrecognizable objects. Thus, evaluating only the first three metrics is unsuitable for this task.
@@
We also include inpainting results obtained by repeating the object generation for each mask for completeness. These alternatives benefit from untruncated prompts and precise knowledge of the inpainting locations, giving them an advantage over other approaches. Despite this, our results show that the RCA approach achieves comparable outcomes while significantly improving efficiency by completing the inpainting process in a single step. Further analysis of generation across multiple mask settings is in the supplementary material.
Finally, we evaluated the entire pipeline using our fine-tuned Stable Diffusion with RCA and generated prompts. As expected, this model achieves the second-best FID score and outperforms SD-2-Inp-HQPrompt in inpainting ground truth entities, showing significant improvement even without access to ground truth prompts.
The improved region-prompt alignment from RCA is confirmed by results on DCI (Table 4), where the RCA variant of the model outperforms the alternatives in CLIPSim.
@@ A Additional Dataset Information
This section provides additional details on the automatically annotated dataset of digitized images of artworks from $\mathcal{A}rt\mathcal{G}raph$. As described in the main paper, the images were downloaded as thumbnails from the WikiArt API, using the formats HalfHD, Large, and PinterestLarge. For each image, we selected the highest resolution available. The counts and resolutions of the images are summarized in Table 5.
@@ B Additional Experiments
Figure 8: Comparative analysis on multi-mask inpainting.
@@ B.1 Analysis on the Number of Masks
We compare the base Stable Diffusion-2.0-Inpainting model, a fine-tuned version that repeats inpainting for each mask, and our RCA fine-tuned version, evaluating performance across varying numbers of masks using ground-truth prompts truncated to a maximum of 10 words (Fig. 8).
The figure shows that our method consistently achieves better region-prompt alignment than the base Stable Diffusion model, regardless of the number of masks. Notably, as the number of masks increases, prompt-following scores decline, with the largest drop occurring when transitioning from single-mask to multi-mask inpainting, highlighting the increased difficulty of the task.
Interestingly, RCA slightly outperforms the repeated inpainting variant in the single-mask setting, while performing similarly in settings with more masks. This is significant because the RCA model requires a single backward pass and does not benefit from limiting the inpainting region to a single prompt. Both methods outperform the base Stable Diffusion-2.0-Inpainting model.
We also report the overall CLIPSim-T2I scores based on global image descriptions from Kosmos-2 during annotation. Both fine-tuned variants exceed the baseline, with RCA slightly outperforming the repeated inpainting approach.
@@
Figure 9: Qualitative comparison between repeated mask inpainting and multi-mask inpainting with RCA.
A qualitative comparison is shown in Fig. 9. Analyzing the generated images reveals that while repeating the inpainting process for each mask improves region-prompt alignment, it negatively impacts the overall aesthetics of the images. This approach sometimes introduces objects into the scene without meaningful integration with the surrounding elements. This issue is particularly evident in the second row of examples, where RCA prioritizes creating a cohesive scene by positioning the basket to interact meaningfully with its surroundings, even if this comes at the expense of precise positioning. In contrast, the repeated inpainting method appears to “paste” the basket into the scene without establishing such connections.
Moreover, images produced with repeated inpainting exhibit reduced stylistic coherence and are more prone to unintended modifications in regions that should remain untouched. These observations highlight the limitations of relying solely on numerical metrics to evaluate inpainting models. While such metrics provide a baseline assessment of performance, they often fail to capture these nuanced aspects, which are especially significant in the context of artistic imagery.